This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
    NA61/SHINE logo
Message ID: 11     Entry time: Mon, 18.07.2016 17:21:50
Author: M. Kuich, F. Mercury, E. Kaptur on behalf of the NA61/SHINE collaboration 
Subject: Run settings p+Pb @ 80GeV/c 
Target: Lead 5 mm

Beam: protons
Beam momentum: 80 GeV/c

Beamline settings:
http://na61-elog.cern.ch/Beam+logbook/181

Beam vertical counters position:
V0: 56.9 mm
V1: 82.0 mm
S1: aligned by geometers
S2: 84.5 mm


S4 horizontal posotion: 14 mm to Jura side, detector in the beam with full magnetic field
S5 position: out of the beam (h: -400, v: 16.9)

Thresholds:
V0: -130 mV
V1: -518 mV
S1: -158 mV
S2: -153 mV
S4: -338 mV
S5: -289 mV

Trigger:
There was problem with pattern units copying. Runs 23134 - 23183 can do not have trigger pattern units copied to raw/shoe files (!)

Trigger file in use:
15/07-18/07 : 2016_test_pPb_80Gev.cnf
since 18/07 : config-14688126241557.cnf

Trigger settings:
T1 = S1*S2*V0bar*V1bar*Ced6
T2 = S1*S2*S4bar*V0bar*V1bar*Ced6
T3 = S1*S2*V0bar*V1bar
T4 = S1*S2*S4bar*V0bar*V1bar

Trigger scaler factors:
T1 = 200
T2 = 1
T3 = 4095
T4 = 4095

HV settings:
https://na61-elog.cern.ch/Experts+logbook/263
S4 HV: +2200 V


Magnets ON - field corresponding to beam momentum
Vxt 1 - 2397.7 A
Vxt 2 - 1835.6 A


PSD:
within runs 23134 - 23237 there were 2 PSD trigger scintillators in "beam 0" (not in the beam!!) position
within runs 23142 - 23161 mod#12 has faulty FPGA channels
PSD position:
Z at 20.5m from target
platform X = 660 (5cm shift down)
platform Y = 960 (25cm shift to Jura)
NOTE: During all data taking there was "small module" in front of modules 21 and 22!


TOF:
ToF-R: worked in production runs
ToF-L:
- within runs 23134 - 23197 partially worked (64 dead pixels due to bad TDC)
- within runs 23198 - 23206 partially worked (17 dead pixels due to missing trigger cable connection)
- since 23207 seems to be working
ELOG V3.1.3-7933898